英語 (English)

基盤英語 (Basic English) (工 ((知 A)1 年))

ニムチャックアーレン · 非常勤講師 1 単位前期木 1 ・ 2

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】 Students will have a chance to practice using English. 【授業の概要】Students will discuss a different topic each week, practicing the

English necessary for each week’s lesson. 【キーワード】 Practical English; conversation 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students should come to class and try to use only English, without

fear of making mistakes. Mistakes will happen, and students should not worry when they make them. Enjoy the class, have fun, and use English.

【授業の計画】 1. Introductions

2. Classroom English

3. What did you do last weekend?

4. Which is bigger?

5. Time

6. I like cheese pizza. Me too!

7. First, cut some tomatoes.

8. Mid-Term Test

9. Future plans

10. How often do you exercise?

11. Can you swim?

12. Fast food

13. Holidays and celebrations

14. I Prefer dark chocolate

15. Can you speak any other languages?

16. Final Test 【教科書】無【参考書等】 Practical English Textbook - Basic (\begin{matrix}Y \\ = \end{matrix}500) - available from teacher,

2nd week 【成績評価の方法】 Students’ grades will be as follows: Attendance = 50% of students’ grades. Mid- term test = 25% of students’ grades. Final test = 25% of students’ grades. Final grades will consist of an attendance score and test scores. (ie. 50% + 25% + 25% = 100%)

【再試験の有無】無

【受講者へのメッセージ】 Have fun, enjoy and speak English!

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220860 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ ニムチャック . (オフィスアワー: canadaclubjp@bloom.ocn.ne.jp)

English

Basic English (工 ((知 A)1 年))

ARLEN NIMCHUK · Part-time Lecturer 1 unit 前期木 1 ・ 2

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ Students will have a chance to practice using English. Outline⟩ Students will discuss a different topic each week, practicing the English

necessary for each week’s lesson. �������⟩ Practical English; conversation Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students should come to class and try to use only English, without fear of

making mistakes. Mistakes will happen, and students should not worry when they make them. Enjoy the class, have fun, and use English.

Schedule⟩

1. Introductions

2. Classroom English

3. What did you do last weekend?

4. Which is bigger?

5. Time

6. I like cheese pizza. Me too!

7. First, cut some tomatoes.

8. Mid-Term Test

9. Future plans

10. How often do you exercise?

11. Can you swim?

12. Fast food

13. Holidays and celebrations

14. I Prefer dark chocolate

15. Can you speak any other languages?

16. Final Test Textbook⟩ 無 Reference⟩ Practical English Textbook - Basic (\begin{matrix}Y \\ = \end{matrix}500) - available from teacher,

2nd week Evaluation Criteria⟩ Students’ grades will be as follows: Attendance = 50% of students’ grades. Mid- term test = 25% of students’ grades. Final test = 25% of students’ grades. Final grades will consist of an attendance score and test scores. (ie. 50% + 25% + 25% = 100%)

Re-evaluation⟩ 無

Message⟩ Have fun, enjoy and speak English!

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220860 Contact (Office-Hour, Room, E-mail)⟩

⇒ NIMCHUK . (Office Hour: canadaclubjp@bloom.ocn.ne.jp)